4. You can insure a $50,000 diamond for its total value by paying a premium of D dollars. If

the probability of theft in a given year is estimated to be 0.01, what premium should the

insurance company charge if it wants the expected gain to equal $1000?


1
Expert's answer
2022-03-15T09:29:16-0400


Probability Cost to the company

0.99 D

0.01. -50000

The expected value is:

E = 0.99(D) + 0.01(-50000)

Given that the expected gain is $1000

Then:

0.99(D) + 0.01(-50000) = 1000

0.99D = 1000 + 500

D = 1500/0.99

D= 1515.15


Answer: $1515.15